OpenStudy (anonymous):

A advertising poster in the shape of a triangle has one side that is 5 inches more than the length of the shortest side, and another side that is 3 inches less than twice the shortest side. Find the dimensions of the poster if its perimeter is 26 inches.

OpenStudy (anonymous):

Alright. So here's how we do this one:

OpenStudy (anonymous):

The perimeter is 26, right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Sorry! My internet died suddenly. Ok, where was I....

OpenStudy (anonymous):

5*x

OpenStudy (anonymous):

Ah yes... So we basically know that Side 1 + side 2 + side 3 = 26 Sound good so far?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

6 9 and 11? yay! win!

OpenStudy (anonymous):

Okie dokie. We'll let side one be the shortest.

OpenStudy (anonymous):

3-x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@woogy: the answer isn't as important as knowing how to do it. I too can just give him the answers in a few seconds. Explaining it takes longer, but I want him to learn how to do it.

OpenStudy (anonymous):

Ohhh... my bad. Want the explanations then?

OpenStudy (anonymous):

5 inches more than the length of the shortest side So if we have a side that is 5 inches more than the shortest side, how do we show this?

OpenStudy (anonymous):

5+x?

OpenStudy (anonymous):

Excellent! So keep note of that: Side 1 = x Side 2 = x+5

OpenStudy (anonymous):

Now here's the hard part... how do we show this: and another side that is 3 inches less than twice the shortest side

OpenStudy (anonymous):

It might look scary, but you can do it!

OpenStudy (anonymous):

3-x^2

OpenStudy (anonymous):

Close. But x^2 would be like saying: Twice of 4 is 4^2, which is 16 And we know 16 is not twice of 4.

OpenStudy (anonymous):

3-x*2

OpenStudy (anonymous):

How would you say "two times x", in other words.

OpenStudy (anonymous):

Perfect! Or just 3-2x in more conventional terms

OpenStudy (anonymous):

side 3 3-2x

OpenStudy (anonymous):

Alright, so we have: Side 1 is x Side 2 is x+5 Side 3 is 2x-3

OpenStudy (anonymous):

Also, I made a mistake. It should have said 2x-3 not 3 -2x

OpenStudy (anonymous):

is there a difference 2x-3 or 3-2x?

OpenStudy (anonymous):

The reason is because saying 3-2x would mean "it's 3 inches minus twice the size of side 1"

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

x+x+5+2x-3

OpenStudy (anonymous):

Yeah. If you plug in a random number, like say 0, the difference is that 2x-3 would equal -3 and 3-2x would equal positive 3.

OpenStudy (anonymous):

Good! Now what do the three sides = to?

OpenStudy (anonymous):

x+x5+2x-3=26

OpenStudy (anonymous):

The key is being able to write the mthematic phrases down correctly... rest is just additions.. Side A (shortest) = x Side B = x + 5 (add 5 inches to the lenght of the shortest) Side C = 2x-3 (two time the shorest [which is x], minus 3) Total = 26 (perimeter) Put it all in 1 equation... Side A + Side B + Side C = 26 Now replace the words with numbers... x + x+5 + 2x-3 = 26 Put all the x's on one side, and the numbrs on the other side... x + x + 2x = 26 - 5 + 3 Add and simplify... 4x = 24 Isolate the x... To do so we divide by 4 each side, so keep the equation balanced... (4x)/4 = 24/4 x = 6

OpenStudy (anonymous):

Good work woogy.

OpenStudy (anonymous):

Am I going too fast or does it make sense?

OpenStudy (anonymous):

Excellent, boog1! Now just do the finishing touches.

OpenStudy (anonymous):

so 6 is for which side?

OpenStudy (anonymous):

When unsure... go back and look at your mathematical phrases

OpenStudy (anonymous):

We said Side A (shortest) = ? Side B = ? Side C = ? Always keep those written somewhere in case you get lost... It happend ot me so many times also that when I got the answer (like x = 6) and I was like... Umm ok, what now?? :P

OpenStudy (anonymous):

6,11,9

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

ok! thanks for your help guys!! I will be back again!!!

OpenStudy (anonymous):

Sounds great. Hope we made math less scary. It's a beautiful thing when it makes sense.

OpenStudy (anonymous):

If you want a tip... I'm a visual kind of person... It reeealllly helps me to draw and write the key infos when solving problems... Like dont be scared to draw the triangle, and write x on a side,x+5 on the other and 2x-3 on the other... no matther how you shape the triangle, its just to help keep things concrete in your head

OpenStudy (anonymous):

i used to like math till I started algebra

OpenStudy (anonymous):

hahaha ;) I was horrible at algebra because i didn't understand why X was a different thing in each problem... yet we always write x :P Until a teach said "you have to see X like an empty box... if you're confused, instead of saying X, say "box" "... Then it made things conrete for me

OpenStudy (anonymous):

any more problems to solve? maybe the next one you can give it a shot by yourself and we can tell you if you're headed the good/wrong way?

OpenStudy (anonymous):

^ I like that! My problem was remembering when we subtract or when we divide when making an equation. I used to be like: "Given rt = d, how do you solve for r? Me: Well... subtract r from both sides, and you have t = d-r" Lol

OpenStudy (anonymous):

haha, yea... that gets tricky too when priority of equations come into play with parentheses and whatnot... thats why (as done above) i always try to solve equations in 2 steps... first the plus' and minus (+, -) which are easy, then the dividing/multiplying (x, /) once everything is on the right side

OpenStudy (anonymous):

Yep. If only I could revisit "highschool me" and tell him this stuff. Gosh, I would have been an Einstein by now. :P

OpenStudy (anonymous):

hahahahah!! damn right... I'd nail high school!! :P

OpenStudy (anonymous):

Alright, time to find more people to help.. Sorry if i was going too fast earlier, its my 1st time on this site

OpenStudy (anonymous):

You did awesome! Take care :)

OpenStudy (anonymous):

Nah, I found that you taught it pretty well. You provided an explanation, which is perfect. My method only works live, so while I prefer it to just giving an explanation, it's worthless if the person refuses to answer me. Luckily boog1 actually replies to questions :D So yeah, your method's perfect. Though if the person's around, you should try to engage them to see where they make mistakes and explain it to them. It might be obvious to the person that "2 times x is 2x" when you tell it to them, but some folks might accidentally think it's "2 + x" or "x times x" if you didn't tell them.

OpenStudy (anonymous):

Right right :) gonna work on that :)